                      Test of epsf.tex

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be centered:

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be left-adjusted:

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be right-adjusted:

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be centered with abutting arrows on the baseline:

⇒  ⇐

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be left-adjusted with abutting arrows on the baseline, the leftmost of which is in the left margin on the baseline:

⇒  ⇐

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be right-adjusted with abutting arrows on the baseline, the rightmost of which is in the right margin on the baseline:

⇒  ⇐

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

                      Test of epsf.tex

with

 \epsfframetrue and \epsfshowfalse and \epsfshowfilenametrue

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be centered:

                          tepsf.eps

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be left-adjusted:

      tepsf.eps

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be right-adjusted:

                                              tepsf.eps

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be centered with abutting arrows on the baseline:

              ⇒  tepsf.eps  ⇐

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be left-adjusted with abutting arrows on the baseline, the leftmost of which is in the left margin on the baseline:

⇒  tepsf.eps  ⇐

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah

This figure should be right-adjusted with abutting arrows on the baseline, the rightmost of which is in the right margin on the baseline:

                                  ⇒  tepsf.eps  ⇐

Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah. Blah blah blah blah blah blah